Need help! I asked earlier but have yet to get a response and I am trying to study for my exam on Monday. Please help.
1) -2* (-5) - (2 - 2^3)^2 (I got -26)
2) -2(-6t+z) - 4(-5z+6t) (I got -12t+18z)
3) -3(w+7) = 3w +33 (I got -9)
4) 4x-5y=12 6x-7y=18 (need help)
5) graph line y=-x+5 (help please)
6)y+3y+2 / 6y^ -6
Thank you

Answer by rfer(16322)

I looked at the first three, they are correct.
5)
y=-x+5
y=mx+b
slope=m=-1
y int=b=5, (0,5)
negetive slope 1 over down 1 starting at (0,5)